# An Improved Algorithm for Adversarial Linear Contextual Bandits via Reduction

### v1

@@ Abstract
We present an efficient algorithm for linear contextual bandits with adversarial losses and stochastic action sets. Our approach reduces this setting to misspecification-robust adversarial linear bandits with fixed action sets. Without knowledge of the context distribution or access to a context simulator, the algorithm achieves $\widetilde{\mathcal{O}}(\min\{d^{2}\sqrt{T},\sqrt{d^{3}T\log K}\})$ regret and runs in $\mathop{\mbox{poly}}(d,C,T)$ time, where $d$ is the feature dimension, $C$ is an upper bound on the number of linear constraints defining the action set in each round, $K$ is an upper bound on the number of actions in each round, and $T$ is number of rounds. This resolves the open question by Liu et al., (2023) on whether one can obtain $\mathop{\mbox{poly}}(d)\sqrt{T}$ regret in polynomial time independent of the number of actions. For the important class of combinatorial bandits with adversarial losses and stochastic action sets where the action sets can be described by a polynomial number of linear constraints, our algorithm is the first to achieve $\mathop{\mbox{poly}}(d)\sqrt{T}$ regret in polynomial time, while no prior algorithm achieves even $o(T)$ regret in polynomial time to our knowledge. When a simulator is available, the regret bound can be improved to $\widetilde{\mathcal{O}}(d\sqrt{L^{\star}})$, where $L^{\star}$ is the cumulative loss of the best policy.
@@ 1 Introduction
This framework is applicable in settings such as healthcare and recommendation systems, where decisions must be made conditional on context. Prior work on contextual bandits has studied a variety of assumptions on how losses and contexts are generated. While much of the literature assumes i.i.d. losses and arbitrarily chosen action sets (for which a well-known algorithm is LinUCB Li et al., (2010)), we focus on the complementary regime: the action sets are drawn i.i.d. from a fixed distribution $\mathcal{D}$, while the losses may be chosen adversarially.
@@
A first computationally efficient algorithm for this setting was proposed by Neu and Olkhovskaya, (2020) under the assumption that the context (i.e., action set) distribution is known. Since an action set is a subset of $\mathbb{R}^{d}$ (i.e., it lies in the space $2^{\mathbb{R}^{d}}$), the distribution over action sets is in the space $\Delta(2^{\mathbb{R}^{d}})$, which is generally intractable to represent efficiently. This assumption was later removed by subsequent efficient algorithms (Luo et al.,, 2021; Sherman et al.,, 2023; Dai et al.,, 2023; Liu et al.,, 2023). In the setting where the learner has access to a simulator that can generate free contexts (i.e., the learner is able to sample contexts from $\mathcal{D}$ as many times as they want without incurring cost), Dai et al., (2023) shows that a near-optimal regret bound of $\widetilde{\mathcal{O}}(\min\{d\sqrt{T},\sqrt{dT\log K}\})$ is achievable. When the learner has neither knowledge of the context distribution nor simulator access to random context samples, the best known results are by Liu et al., (2023): they provide an algorithm with near-optimal regret $\widetilde{\mathcal{O}}(d\sqrt{T})$ with run time $T^{\Omega(d)}$ and another algorithm with regret $\widetilde{\mathcal{O}}(d^{2}\sqrt{T})$ with run time $\mathop{\mbox{poly}}(d,K,T)$ where $K=\max_{t}|\mathcal{A}_{t}|$. Notably, while the regret bound of this last algorithm is independent of the number of actions $K$, its computational complexity scales polynomially in $K$. In fact, this is the case for all previous algorithms as well (Luo et al.,, 2021; Sherman et al.,, 2023; Dai et al.,, 2023; Liu et al.,, 2023; Liu et al., 2024b, ). This makes them unsuitable for many important combinatorial problems (e.g., $m$-set, shortest path, flow, bipartite matching), where $K$ is usually exponentially large in the dimension $d$.
@@
Table 1: Comparison with state-of-the-art results in adversarial linear contextual bandits. $d$ is the feature dimension, $K$ is an upper bound on the number of actions, and $C$ is an upper bound on the number of linear constraints to describe the convex hull of each action set. It holds that $C\leq K+1$ and in many combinatorial problems we have $C=\mathop{\mbox{poly}}(d)$ and $K=2^{\Omega(d)}$. The run time of the linear optimization oracle of Neu and Valko, (2014) is bounded by $\mathop{\mbox{poly}}(d,C)$ but could also be smaller.
Algorithm Regret (omitting $\log(dT)$ factors) Computation Simulator Feedback
Schneider and Zimmert, (2023)
$\sqrt{dT}$ for the special case
$\mathcal{A}_{t}\subseteq\{\mathbf{e}_{1},\ldots,\mathbf{e}_{d}\}$
@@
$\mathop{\mbox{poly}}(d,T)$ no bandit
Neu and Valko, (2014) $(dT)^{2/3}$
$\mathop{\mbox{poly}}(d,T)$ plus
$T$ oracle calls
no semi-bandit
Dai et al., (2023) $\min\{d\sqrt{T},\sqrt{dT\log K}\}$ $\mathop{\mbox{poly}}(d,K,T)$ yes bandit
Liu et al., (2023) $d\sqrt{T}$ $K\cdot T^{\Omega(d)}$ no bandit
Liu et al., (2023) $d^{2}\sqrt{T}$ $\mathop{\mbox{poly}}(d,K,T)$ no bandit
Ours $\min\{d^{2}\sqrt{T},\sqrt{d^{3}T\log K}\}$
$\mathop{\mbox{poly}}(d,C,T)$
no bandit
Ours $d\sqrt{L^{\star}}$
$\mathop{\mbox{poly}}(d,C,T)$
yes bandit
@@
Our work gives the first algorithm whose computational complexity does not explicitly scale with the number of actions, making adversarial linear contextual bandits applicable to a much wider range of problems. Without simulator access, our method achieves regret $\widetilde{\mathcal{O}}(\min\{d^{2}\sqrt{T},\sqrt{d^{3}T\log K}\})$ and with simulator access this can be improved to $\widetilde{\mathcal{O}}(d\sqrt{L^{\star}})$, where $L^{\star}=O(T)$ is the cumulative loss of the best policy. Our algorithm runs in $O(d,C,T)$ time, where $C$ is the number of linear constraints to describe the convex hull of each action set. Notice that $C\leq K+1$ in general, as the convex hull of an action set of size $K$ can be written as a linear program with at most $K+1$ constraints. On the other hand, in many combinatorial problems, $C=\mathop{\mbox{poly}}(d)$ while $K=2^{\Omega(d)}$. For example, in a shortest path problem with $d$ edges, the set of all paths can be described by a linear program with $O(d)$ constraints, while the number of paths could be of order $2^{\Omega(d)}$. For combinatorial problems with stochastic action sets and adversarial losses, we are only aware of Neu and Valko, (2014) who studied the case where the learner has semi-bandit feedback. Their algorithm achieves $\widetilde{\mathcal{O}}\big{(}(dT)^{2/3}\big{)}$ regret with one call to the linear optimization oracle of the action set per round. Compared to their work, our work weakens the assumption on the feedback (from semi-bandits to full-bandits) and improves the regret bound, but our method is computationally heavier—there are action sets where linear optimization can be solved in polynomial time while having an exponentially large number of constraints, such as spanning trees. How to further improve our computational complexity to match theirs is left as future work. It also remains open how to efficiently achieve the near-optimal bound $\widetilde{\mathcal{O}}(\min\{d\sqrt{T},\sqrt{dT\log K}\})$ without simulators, even when the learner is willing to pay $\mathop{\mbox{poly}}(d,K,T)$ computation. This has only been resolved in the special case where action sets are subsets of the standard basis $\{\mathbf{e}_{1},\ldots,\mathbf{e}_{d}\}$, which is also known as the sleeping bandits problem (Schneider and Zimmert,, 2023).
@@
Our result is achieved by establishing a novel and computationally efficient reduction from adversarial linear contextual bandits to adversarial linear bandits with misspecification. Linear bandits can be viewed as linear contextual bandits with a fixed context and is therefore less challenging than linear contextual bandits. This reduction scheme is related to the work of Hanna et al., (2023), which reduces stochastic linear contextual bandits (with both stochastic contexts and stochastic losses) to misspecified stochastic linear bandits. Compared to Hanna et al., (2023), we make two key advances: First, our reduction is polynomial-time, and we also provide a polynomial-time algorithm for the problem we reduce to, whereas Hanna et al., (2023) focuses solely on the regret bound and it remains open how their reduction can be implemented in polynomial time. Second, we study the strictly more general setting with adversarial losses. A more detailed comparison with Hanna et al., (2023) is provided in Appendix A.
@@
Finally, regarding results in terms of $L^{\star}$, we remark that our bound of $\widetilde{\mathcal{O}}(d\sqrt{L^{\star}})$ can be specialized to the setting of Olkhovskaya et al., (2023) (a slightly different formulation of adversarial contextual bandits). We then obtain a slightly worse rate but, importantly, remove their unrealistic assumption that the context distribution is log-concave.
@@ 2.1 Linear Contextual Bandits
For simplicity, we consider an oblivious adversary. [^1] There are two kinds of results we can get: 1) regret compared to the best policy in the full policy set $\Pi$ under oblivious adversary, and 2) regret compared to the best policy in the linear policy set $\Pi_{\textrm{lin}}$ under adaptive adversary. To achieve 2), it suffices to use misspecification-robust linear bandit algorithms with high-probability bounds in our reduction. This can be achieved by standard techniques of getting high-probability bounds (Lee et al.,, 2020; Zimmert and Lattimore,, 2022). To simplify the exposition, we only focus on the first case. Before any interaction with the learner, the adversary secretly chooses $T$ loss vectors $(\theta_{t})_{t\in[T]}$ where $\theta_{t}\in\mathbb{R}^{d}$ for all $t$. For each round $t=1,2,\ldots,T$, an action set $\mathcal{A}_{t}\subset\mathbb{R}^{d}$ is drawn according to $\mathcal{A}_{t}\overset{\text{i.i.d.}}{\sim}\mathcal{D}$ and revealed to the learner. The learner then chooses an action $a_{t}\in\mathcal{A}_{t}$ and observes loss $\ell_{t}\in[0,1]$ with $\mathbb{E}_{t}[\ell_{t}]=a_{t}^{\top}\theta_{t}$, where $\mathbb{E}_{t}$ denotes the expectation conditioned on the history up to time t1t-1 and $a_{t}$. The expected regret against a fixed policy $\pi\in\Pi$ is defined as
$\mathrm{Reg}_{T}(\pi)=\operatorname*{\mathbb{E}}\left[\sum_{t=1}^{T}\left\langle a_{t}-\pi(\mathcal{A}_{t}),\theta_{t}\right\rangle\right]$
@@ 3.2 Connection between Linear Contextual Bandits and Linear Bandits
1 Input: An adversarial linear bandit algorithm $\mathsf{ALG}$ and a set $\mathcal{S}_{N}$ of NN action sets drawn from $\mathcal{D}$.
@@
$\left\langle\tilde{\psi},\phi\right\rangle$ $=\mathbb{E}_{\mathcal{A}\sim\mathcal{D}}\left[\left\langle\pi_{\phi}(\mathcal{A}),\phi\right\rangle\right]$
$=\mathbb{E}_{\mathcal{A}\sim\mathcal{D}}\left[\min_{a\in\mathcal{A}}\left(a^{\top}\phi\right)\right]$
$\leq\mathbb{E}_{\mathcal{A}\sim\mathcal{D}}\left[\pi(\mathcal{A})^{\top}\phi\right]=\left\langle\psi,\phi\right\rangle.$
@@
Based on Lemma 2, we have the following strategy to execute $y_{t}$. First, decompose $y_{t}$ as a convex combination of extreme points of $\Omega$, i.e., decompose $y_{t}$ as $\sum_{i}\alpha_{i}\psi_{i}$ where $\psi_{1},\psi_{2},\ldots$ are vertices of $\Omega$, and $(\alpha_{1},\alpha_{2}\ldots)$ is a distribution over them. Then for each $\psi_{i}$, find the corresponding $\phi_{i}\in-\mathcal{N}(\Omega,\psi)$ according to Lemma 2. Finally, let $\pi_{t}$ be the policy that mix $\pi_{\phi_{1}},\pi_{\phi_{2}},\ldots$ with weights $(\alpha_{1},\alpha_{2},\cdots)$. This way, we have by Lemma 2,
@@ 3.3 Bounding the error due to the discrepancy between $\widehat{\Omega}$ and $\Omega$
Real regret on $\Omega$ that we care about: $\qquad\mathbb{E}_{\mathcal{A}\sim\mathcal{D}}[\left\langle\pi_{t}(\mathcal{A}),\theta_{t}\right\rangle]-\mathbb{E}_{\mathcal{A}\sim\mathcal{D}}[\left\langle\pi(\mathcal{A}),\theta_{t}\right\rangle],$
Regret on $\widehat{\Omega}$ that the robust linear bandit algorithm can bound: $\qquad\mathbb{E}_{\mathcal{A}\sim\widetilde{\mathcal{D}}}[\left\langle\pi_{t}(\mathcal{A}),\theta_{t}\right\rangle]-\mathbb{E}_{\mathcal{A}\sim\widetilde{\mathcal{D}}}[\left\langle\pi(\mathcal{A}),\theta_{t}\right\rangle].$

### v2

@@ Abstract
We present an efficient algorithm for linear contextual bandits with adversarial losses and stochastic action sets. Our approach reduces this setting to misspecification-robust adversarial linear bandits with fixed action sets. Without knowledge of the context distribution or access to a context simulator, the algorithm achieves $\widetilde{\mathcal{O}}(\min\{d^{2}\sqrt{T},\sqrt{d^{3}T\log K}\})$ regret and runs in $\mathop{\mbox{poly}}(d,C,T)$ time, where $d$ is the feature dimension, $C$ is an upper bound on the number of linear constraints defining the action set in each round, $K$ is an upper bound on the number of actions in each round, and $T$ is number of rounds. This resolves the open question by liu2024bypassing on whether one can obtain $\mathop{\mbox{poly}}(d)\sqrt{T}$ regret in polynomial time independent of the number of actions. For the important class of combinatorial bandits with adversarial losses and stochastic action sets where the action sets can be described by a polynomial number of linear constraints, our algorithm is the first to achieve $\mathop{\mbox{poly}}(d)\sqrt{T}$ regret in polynomial time, while no prior algorithm achieves even $o(T)$ regret in polynomial time to our knowledge. When a simulator is available, the regret bound can be improved to $\widetilde{\mathcal{O}}(d\sqrt{L^{\star}})$, where $L^{\star}$ is the cumulative loss of the best policy.
@@ 1 Introduction
This framework is applicable in settings such as healthcare and recommendation systems, where decisions must be made conditional on context. Prior work on contextual bandits has studied a variety of assumptions on how losses and contexts are generated. While much of the literature assumes i.i.d. losses and arbitrarily chosen action sets (for which a well-known algorithm is LinUCB (li2010contextual)), we focus on the complementary regime: the action sets are drawn i.i.d. from a fixed distribution $\mathcal{D}$, while the losses may be chosen adversarially.
@@
A first computationally efficient algorithm for this setting was proposed by neu20 under the assumption that the context (i.e., action set) distribution is known. Since an action set is a subset of $\mathbb{R}^{d}$ (i.e., it lies in the space $2^{\mathbb{R}^{d}}$), the distribution over action sets is in the space $\Delta(2^{\mathbb{R}^{d}})$, which is generally intractable to represent efficiently. This assumption was later removed by subsequent efficient algorithms (luo2021policy; sherman2023improved; dai2023refined; liu2024bypassing). In the setting where the learner has access to a simulator that can generate free contexts (i.e., the learner is able to sample contexts from $\mathcal{D}$ as many times as they want without incurring cost), dai2023refined shows that a near-optimal regret bound of $\widetilde{\mathcal{O}}(\min\{d\sqrt{T},\sqrt{dT\log K}\})$ is achievable, where $K=\max_{t}|\mathcal{A}_{t}|$. When the learner has neither knowledge of the context distribution nor simulator access to random context samples, the best known results are by liu2024bypassing: they provide an algorithm with near-optimal regret $\widetilde{\mathcal{O}}(d\sqrt{T})$ with run time $T^{\Omega(d)}$ and another algorithm with regret $\widetilde{\mathcal{O}}(d^{2}\sqrt{T})$ with run time $\mathop{\mbox{poly}}(d,K,T)$. Notably, while the regret bound of this last algorithm is independent of the number of actions $K$, its computational complexity scales polynomially in $K$. In fact, this is the case for all previous algorithms as well (luo2021policy; sherman2023improved; dai2023refined; liu2024bypassing; liu2024towards). This makes them unsuitable for many important combinatorial problems (e.g., $m$-set, shortest path, flow, bipartite matching), where $K$ is usually exponentially large in the dimension $d$.
@@
Table 1: Comparison with state-of-the-art results in adversarial linear contextual bandits. $d$ is the feature dimension, $K$ is an upper bound on the number of actions, and $C$ is an upper bound on the number of linear constraints to describe the convex hull of each action set. It holds that $C\leq K+1$ and in many combinatorial problems we have $C=\mathop{\mbox{poly}}(d)$ and $K=2^{\Omega(d)}$. The run time of the linear optimization oracle of neu2014online is bounded by $\mathop{\mbox{poly}}(d,C)$ but could also be smaller.
Algorithm Regret Computation Simulator Feedback
(omitting $\log(dT)$ factors)
schneider2023optimal $\sqrt{dT}$ for the special case $\mathcal{A}_{t}\subseteq\{\mathbf{e}_{1},\ldots,\mathbf{e}_{d}\}$ $\mathop{\mbox{poly}}(d,T)$ no bandit
@@
neu2014online $(dT)^{2/3}$ $\mathop{\mbox{poly}}(d,T)$ plus $T$ oracle calls no semi-bandit
dai2023refined $\min\{d\sqrt{T},\sqrt{dT\log K}\}$ $\mathop{\mbox{poly}}(d,K,T)$ yes bandit
liu2024bypassing $d\sqrt{T}$ $K\cdot T^{\Omega(d)}$ no bandit
liu2024bypassing $d^{2}\sqrt{T}$ $\mathop{\mbox{poly}}(d,K,T)$ no bandit
Ours $\min\{d^{2}\sqrt{T},\sqrt{d^{3}T\log K}\}$ $\mathop{\mbox{poly}}(d,C,T)$ no bandit
Ours $d\sqrt{L^{\star}}$ $\mathop{\mbox{poly}}(d,C,T)$ yes bandit
@@
Our work gives the first algorithm whose computational complexity does not explicitly scale with the number of actions, making adversarial linear contextual bandits applicable to a much wider range of problems. Without simulator access, our method achieves regret $\widetilde{\mathcal{O}}(\min\{d^{2}\sqrt{T},\sqrt{d^{3}T\log K}\})$ and with simulator access this can be improved to $\widetilde{\mathcal{O}}(d\sqrt{L^{\star}})$, where $L^{\star}=O(T)$ is the cumulative loss of the best policy. Our algorithm runs in $O(d,C,T)$ time, where $C$ is the number of linear constraints to describe the convex hull of each action set. Notice that $C\leq K+1$ in general, as the convex hull of an action set of size $K$ can be written as a linear program with at most $K+1$ constraints. On the other hand, in many combinatorial problems, $C=\mathop{\mbox{poly}}(d)$ while $K=2^{\Omega(d)}$. For example, in a shortest path problem with $d$ edges, the set of all paths can be described by a linear program with $O(d)$ constraints, while the number of paths could be of order $2^{\Omega(d)}$. For combinatorial problems with stochastic action sets and adversarial losses, we are only aware of neu2014online who studied the case where the learner has semi-bandit feedback. Their algorithm achieves $\widetilde{\mathcal{O}}\big((dT)^{2/3}\big)$ regret with one call to the linear optimization oracle of the action set per round. Compared to their work, our work weakens the assumption on the feedback (from semi-bandits to full-bandits) and improves the regret bound, but our method is computationally heavier—there are action sets where linear optimization can be solved in polynomial time while having an exponentially large number of constraints, such as spanning trees. How to further improve our computational complexity to match theirs is left as future work. It also remains open how to efficiently achieve the near-optimal bound $\widetilde{\mathcal{O}}(\min\{d\sqrt{T},\sqrt{dT\log K}\})$ without simulators, even when the learner is willing to pay $\mathop{\mbox{poly}}(d,K,T)$ computation. This has only been resolved in the special case where action sets are subsets of the standard basis $\{\mathbf{e}_{1},\ldots,\mathbf{e}_{d}\}$, which is also known as the sleeping bandits problem (schneider2023optimal).
@@
Our result is achieved by establishing a novel and computationally efficient reduction from adversarial linear contextual bandits to adversarial linear bandits with misspecification. Linear bandits can be viewed as linear contextual bandits with a fixed context and is therefore less challenging than linear contextual bandits. This reduction scheme is related to the work of hanna2023contexts, which reduces stochastic linear contextual bandits (with both stochastic contexts and stochastic losses) to misspecified stochastic linear bandits. Compared to hanna2023contexts, we make two key advances: First, our reduction is polynomial-time, and we also provide a polynomial-time algorithm for the problem we reduce to, whereas hanna2023contexts focus solely on the regret bound and it remains open how their reduction can be implemented in polynomial time. Second, we study the strictly more general setting with adversarial losses. A more detailed comparison with hanna2023contexts is provided in Appendix A.
@@
Finally, regarding results in terms of $L^{\star}$, we remark that our bound of $\widetilde{\mathcal{O}}(d\sqrt{L^{\star}})$ can be specialized to the setting of olkhovskaya2024first (a slightly different formulation of adversarial contextual bandits). We then obtain a slightly worse rate but, importantly, remove their unrealistic assumption that the context distribution is log-concave.
@@ 2.1 Linear Contextual Bandits
For simplicity, we consider an oblivious adversary. [^1] There are two kinds of results we can get: 1) regret compared to the best policy in the full policy set $\Pi$ under oblivious adversary, and 2) regret compared to the best policy in the linear policy set $\Pi_{\textrm{lin}}$ under adaptive adversary. To achieve 2), it suffices to use misspecification-robust linear bandit algorithms with high-probability bounds in our reduction. This can be achieved by standard techniques of getting high-probability bounds (lee2020bias; zimmert2022return). To simplify the exposition, we only focus on the first case. Before any interaction with the learner, the adversary secretly chooses $T$ loss vectors $(\theta_{t})_{t\in[T]}$ where $\theta_{t}\in\mathbb{R}^{d}$ for all $t$. For each round $t=1,2,\ldots,T$, an action set $\mathcal{A}_{t}\subset\mathbb{R}^{d}$ is drawn according to $\mathcal{A}_{t}\overset{\text{i.i.d.}}{\sim}\mathcal{D}$ and revealed to the learner. The learner then chooses an action $a_{t}\in\mathcal{A}_{t}$ and observes loss $\ell_{t}\in[0,1]$ with $\mathbb{E}_{t}[\ell_{t}]=a_{t}^{\top}\theta_{t}$, where $\mathbb{E}_{t}$ denotes the expectation conditioned on the history up to time t1t-1 and $a_{t}$. The expected regret against a fixed policy $\pi\in\Pi$ is defined as
$\mathrm{Reg}_{T}(\pi)=\operatorname*{\mathbb{E}}\left[\sum_{t=1}^{T}\left\langle a_{t}-\pi(\mathcal{A}_{t}),\theta_{t}\right\rangle\right],$
@@ 3.2 Connection between Linear Contextual Bandits and Linear Bandits
1Input: An adversarial linear bandit algorithm $\mathsf{ALG}$ and a set $\mathcal{S}_{N}$ of NN action sets drawn from $\mathcal{D}$.
@@
$\left\langle\tilde{\psi},\phi\right\rangle$ $=\mathbb{E}_{\mathcal{A}\sim\mathcal{D}}\left[\left\langle\pi_{\phi}(\mathcal{A}),\phi\right\rangle\right]=\mathbb{E}_{\mathcal{A}\sim\mathcal{D}}\left[\min_{a\in\mathcal{A}}\left(a^{\top}\phi\right)\right]\leq\mathbb{E}_{\mathcal{A}\sim\mathcal{D}}\left[\pi(\mathcal{A})^{\top}\phi\right]=\left\langle\psi,\phi\right\rangle.$
@@
Based on Lemma 2, we have the following strategy to execute $y_{t}$. First, decompose $y_{t}$ as a convex combination of extreme points of $\Omega$, i.e., decompose $y_{t}$ as $\sum_{i}\alpha_{i}\psi_{i}$ where $\psi_{1},\psi_{2},\ldots$ are vertices of $\Omega$, and $(\alpha_{1},\alpha_{2}\ldots)$ is a distribution over them. Then for each $\psi_{i}$, find the corresponding $\phi_{i}\in-\mathcal{N}(\Omega,\psi)$ according to Lemma 2. Finally, let $\pi_{t}$ be the policy that mixes $\pi_{\phi_{1}},\pi_{\phi_{2}},\ldots$ with weights $(\alpha_{1},\alpha_{2},\cdots)$. This way, we have by Lemma 2,
@@ 3.3 Bounding the error due to the discrepancy between $\widehat{\Omega}$ and $\Omega$
Real regret on $\Omega$ that we care about: $\ \ \mathbb{E}_{\mathcal{A}\sim\mathcal{D}}[\left\langle\pi_{t}(\mathcal{A}),\theta_{t}\right\rangle]-\mathbb{E}_{\mathcal{A}\sim\mathcal{D}}[\left\langle\pi(\mathcal{A}),\theta_{t}\right\rangle],$
Regret on $\widehat{\Omega}$ that the robust linear bandit algorithm can bound: $\ \ \mathbb{E}_{\mathcal{A}\sim\widetilde{\mathcal{D}}}[\left\langle\pi_{t}(\mathcal{A}),\theta_{t}\right\rangle]-\mathbb{E}_{\mathcal{A}\sim\widetilde{\mathcal{D}}}[\left\langle\pi(\mathcal{A}),\theta_{t}\right\rangle].$